OpenStudy (anonymous):

what is 40% as a decimal?? what is 40% as a decimal?? @Mathematics

OpenStudy (anonymous):

.4

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

so is 9% 9? or is it 0.09

OpenStudy (anonymous):

9 is .09

OpenStudy (anonymous):

it is 0.09

OpenStudy (anonymous):

I would not call you the "Mathmaster" :P This is pretty basic math, hehehe :D

OpenStudy (anonymous):

im in 8th grade tho...lol

OpenStudy (anonymous):

what about 2.49%

OpenStudy (anonymous):

.0249

OpenStudy (anonymous):

It's whatever%/100

OpenStudy (anonymous):

but still these questions are worth for class 4th student

OpenStudy (anonymous):

14 1/2% what!?!?

OpenStudy (anonymous):

40/100 = 0.40

OpenStudy (anonymous):

14 1/2 / 100 = 29/2 / 100 = 29/2 * 1/100 = 29/200 = 0.145

OpenStudy (anonymous):

ooooohhhh, i see now, thanks!

OpenStudy (anonymous):

how would i figure out 40% of 80??

OpenStudy (anonymous):

(40/100)*80 = ?

OpenStudy (anonymous):

40/100 * 80